THE NEAR ULTRAVIOLET BAND SYSTEM OF SINGLET METHYLENE

\underline{P. R. BUNKER}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; S. N. YURCHENKO, PER JENSEN, YAN LI AND R. J. BUENKER, FB 9 - Theoretische Chemie, Bergische Universität - Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany.

In a classic paper entitled “The spectrum and structure of singlet CH_{2}” by G. Herzberg and J. W. C. Johns (Proc. Roy. Soc. A295, 107-128 (1966)) the analysis of the b̃ ^{1}B_{1}←ã ^{1}A_{1} red absorption band system of CH_{2} is discussed in detail for the first time. In addition to that band system the observation of a fragment of a weak near ultraviolet absorption band system is reported. The three observed bands of the system could not be vibrationally assigned or rotationally analyzed but it was pointed out that they probably involve absorption into the second excited singlet state, c̃^{1}A_{1}. We show this supposition to be true here by simulation. In order to simulate the spectrum we have calculated ab initio the c̃-ã and c̃-b̃ transition moment surfaces, and used the MORBID and RENNER program systems with previously determined potential energy surfaces for the ã, b̃ and c̃ states in a calculation of the energy levels and wavefunctions. We find that the three bands seen by Herzberg and Johns are part of the c̃←(ã/b̃) system but that all of the bands of the system above about 31 000 cm^{−1} are missing as a result of c̃ state predissociation. We can vibrationally assign the bands, but the weakness of the spectrum, and the presence of perturbations, make it impossible for us to assign all the rotational structure. It would be worthwhile to record this spectrum again using a more sensitive modern laser based technique in order to probe singlet state CH_{2} at these energies.